please help me solve this question. the measure of the supp. of an angle exceeds twice the measure of the comp. of the angle by 20. find the measure of 1/2 the comp.
is this how it is set up? 180-x= 2(90-x)+20
please help...

Answer by rapaljer(4671)

You set it up right. Now solve for x.

180-x=2(90-x)+20
180 - x = 180 -2x + 20

Subtract 180 from each side:
-x=-2x+20

Add +2x to each side:
-x + 2x = -2x + 2x + 20
x= 20

Now, answer the question. Find one half of the complement of the angle.

The angle was 20 degrees, so the complement of this angle would be 70. Now, half of 70 is 35 degrees. Does that look right?

R^2

Dr. Robert J. Rapalje